OpenStudy (anonymous):

Find the height of a right circular cone whose radius is 5 cm and volume is 100π cm3.

OpenStudy (anonymous):

Hint: the volume of a cone is given by: \[V=\frac{ 1 }{ 3 }\pi r^2 h\] the radius is given.

OpenStudy (anonymous):

Jim Debt was reviewing the total accounts receivable. This month he received $80,000 from Credit customers. This represented 40 percent of all receivables due. The total amount of accounts receivable is:

OpenStudy (anonymous):

\[100\pi=\frac{ 1 }{ 3 }5^{2}*\pi*h\]

OpenStudy (anonymous):

Thanks!(:

OpenStudy (anonymous):

You're welcome :)